3. Discharge of the potential across a charged capacitor. A 330 uF (microfarad) capacitor is charged to 1 volt and then discharged through a 100kS2 (kilo ohm) resistor, and the following data were obtained (the voltage was measured at 5 second intervals) Page 3 of 4 TimeVoltage sec) (millivolts(sec) (millivolts TimeVoltage 50 830 743 656 580 513 453 401 350 310 60 65 70 75 80 85 90 269 243 215 190 167 150 131 114 103 25 30 35 40 45 The data can be fit to the equation Plot on a semi-log graph paper and determine the time constant -R
Answer #1

When voltage and time are plotted on semi-log graph paper, the slope of line = -(1/RC)

The slope of line from plotted data is -0.0249

this means, -(1/RC) = -0.0249

ot 1 / RC = 0.0249

or RC = 1 / 0.0249

or RC = 40.16

Hence time constant \tau = RC = 40.16